# In-Context Freeze-Thaw Bayesian Optimization for Hyperparameter Optimization

### v1

@@ 4.1 Dynamic Surrogate Model (FT-PFN)
Further details about meta-training FT-PFN can be found in Appendix A, including the basis curves, their parameters, and illustrations of samples from our learning curve prior (Appendix A.1); the specific procedure we use for generating our meta-training data (Appendix A.2); and the architecture and hyperparameters used (A.3).
4.2 Dynamic Acquisition Function (MFPI-random)
Following the Freeze-Thaw Bayesian Optimization framework (Section 3.2), we continue training the configuration $\lambda\in\Lambda$ that maximizes some acquisition function. Here, we assume a general multi-fidelity (MF) acquisition function (AF) represented as,
$\texttt{MF-AF}(\lambda,z~|~f_{\text{threshold}}):\Lambda\times~\mathcal{Z}\rightarrow\mathbb{R}\text{,}$ (4)
where AF is a general acquisition function, $\lambda\in\Lambda$ a hyperparameter configuration, $z\in\mathcal{Z}$ the fidelity (e.g. number of training steps), and $f_{\text{threshold}}$ a function of $f_{\text{best}}$ (best observed performance). For example, fixing $z=b_{\text{max}}$ and $f_{\text{threshold}}=f_{\text{best}}$, recovers a vanilla AF such as Expected Improvement (EI) or Probability of Improvement (PI) (Mockus et al., 1978).
The quality of uncertainty calibration in surrogate predictions depends on the task, surrogate training set quality, and extrapolation length (horizon). Meanwhile, the AF score is calibrated by an improvement threshold. In our work, we choose PI as the base AF as it is less likely to drive search towards high uncertainty regions along the hyperparameter and fidelity subspaces. Instead of using a fixed threshold or a fixed extrapolation horizon for this PI calculation, we explore a range of possible thresholds and horizons by random sampling from either fixed or problem-specific bounds. Such sampling procedure is undertaken every iteration and is akin to an acquisition selection from a portfolio of different PIs. In contrast to integrating over all thresholds (leading to Expected Improvement, EI), this allows for greater diversification in the exploration process. We posit that this hedging with a portfolio of AFs in each iteration benefits freeze-thaw setups where queries are more granular than standard BO. The result is a simple, parameter-free AF with a balanced exploration-exploitation trade-off,
@@ 4.2 Dynamic Acquisition Function (MFPI-random)
$\texttt{MFPI-random}(\lambda,z^{\text{rand}}~|~f_{\text{threshold}}^{\text{rand}}):\Lambda\times~\mathcal{Z}\rightarrow\mathbb{R}\text{,}$ (5)
which evaluates to the predicted likelihood that a candidate configuration $\lambda\in\Lambda$ after $z^{\text{rand}}$ more steps of training obtains a model exceeding the $f_{\text{threshold}}^{\text{rand}}$ performance threshold, where $z^{\text{rand}}\sim\mathcal{U}(1,b_{\text{max}})$ is the sampled extrapolation horizon, and $f_{\text{threshold}}^{\text{rand}}=f_{\text{best}}\cdot(1-\tau)$, with $log_{10}(\tau)\sim\mathcal{U}(-4,-1)$ as the sampled scaling for the threshold. Further details, as well as pseudo code (Algorithm 2) can be found in Appendix A.4.
@@ 5 Experiments
Our main baselines are both other recent freeze-thaw approaches: DyHPO (Wistuba et al., 2022) and DPL (Kadra et al., 2023). We use the publicly available implementations for DyHPO and DPL with default hyperparameter values.
@@ 5.1 Cost and Quality of Predictions
In this section, we compare the predictive capabilities of FT-PFN to that of existing surrogate models, including the deep Gaussian process of DyHPO (Wistuba et al., 2022) and the deep ensemble of power laws model of DPL (Kadra et al., 2023). We also consider a variant of FT-PFN trained on the same prior, but not taking the hyperparameters as input (referred to as ”no HPs”). This variant bases its predictions solely on a set of partially observed learning curves.
@@
Figure 3: Comparison of our method against state-of-the-art baselines on all 3 benchmarks. First row shows normalized regret aggregated across multiple tasks in each benchmark (See Appendix B for benchmark details, and the results per task can be found in Appendix E). Second row shows the average ranks of each method.
@@ 5.2 Assessment of HPO performance
In this section, we present an extensive empirical comparison, showing the merits of our method on HPO tasks across a variety of tabular benchmarks (see Appendix B) in the low budget regime. Additionally to DPL and DyHPO, we include Hyperband, ASHA, Gaussian Process-based FT-BO (using DyHPO’s one-step EI acquisition function) and uniform random search, as baselines. Each algorithm is allocated a total budget of 1000 steps for every task, which corresponds to 20 full trainings on LCBench and Taskset  with each task being repeated 10 times, each time starting from a different random configuration (see Algorithm 1). We report two complementary metrics: The normalized regret, capturing performance differences, and the average rank of each method, capturing the relative order. Formally, the normalized regret corresponds to a $[0,1]$ normalization of the observed error w.r.t to the best (lowest) and worst (highest) errors recorded by all algorithms on the task.
@@
Figure 3 presents the comparative results per benchmark family. The results validate the superiority of freeze-thaw approaches (ifBO, DyHPO, and DPL) compared to standard approaches (Hyperband, ASHA, and random search) for low-budget settings. Most notably, these results establish the promise of ifBO, which either outperforms (on LCBench and Taskset) or competes closely (on PD1) with DPL and DyHPO. ifBO is also consistently the best on average rank across all benchmarks (see Appendix C.2, Figure 7). Appendix E (Figures 11-16) offers a closer look at the raw error metrics for each algorithm per task. These detailed results collectively confirm the robustness of ifBO for HPO tasks, showing its ability to compete if not outperform the most competitive baselines.
@@ 5.3 Ablation on Acquisition
In this section, we evaluate how ifBO performs in combination with other acquisition functions and aim to assess to what extent our novel acquisition function described in Section 4.2 contributes to its HPO success. To this end, we consider ifBO variants combining FT-PFN with and comparing performance against the EI-based acquisitions used in prior-art (Wistuba et al., 2022; Kadra et al., 2023), i.e., EI (one step) predicting one step in the future (DyHPO), and EI (max) predicting at the highest budget $b_{max}$ (DPL). We also include their PI counterparts PI (one step) and PI (max), as well as variants of MFPI-random that only vary the prediction horizon, PI (random horizon) with $T=0$, or only vary the threshold, PI (max, random-T) with $h=b_{max}$. Apart from the methods compared, the experimental setup is identical to that in Section 5.2.
@@
We perform comparable ablation studies for DPL and DyHPO, as detailed in Appendix C.3, to demonstrate the benefits of randomizing the horizon and the threshold.
@@ 6 Conclusion
Despite our promising results, we admit that attaining the sample efficiency required to scale up to modern deep learning (e.g., LMM pretraining) remains a challenging endeavor. Future work should attempt to extend our approach to take advantage of additional sources of prior information, e.g., to do in-context meta-learning, leveraging learning curves on related tasks (Ruhkopf et al., 2022); to incorporate user priors (Müller et al., 2023; Mallik et al., 2023); and additional information about the training process (e.g., gradient statistics). An alternative to scaling up is scaling in parallel. Here, we expect our in-context learning approach to reduce the overhead further, as the online learning/refitting stage occurs on the critical path, while in-context learning during prediction can easily be parallelized. Finally, the current FT-PFN model has some limitations. First, it requires the performance metric and each hyperparameter value to be normalized in $[0,1]$; and supports up to 10 hyperparameters. While we believe that this is reasonable, future work building systems should push these limits, train larger models, on more data, and explore ways to scale to larger context sizes. In summary, there is a lot left unexplored, and we hope that the relative simplicity, efficiency, and public availability of our method, lowers the threshold for future research in this direction.
@@ Acknowledgements
We thank Johannes Hog for his feedback on the earlier version of the paper. We acknowledge funding by the European Union (via ERC Consolidator Grant Deep Learning 2.0, grant no. 101045765), TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215, the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant numbers INST 39/963-1 FUGG and 417962828. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ A.1 The Learning Curve Prior
$\begin{split}\hat{f}(t,\mathcal{E})\sim\mathcal{N}(\hat{f}_{comb}(t,\mathcal{E}),{\sigma^{2}})\text{\quad with \quad}f_{comb}(t,\mathcal{E})=y_{0}+(y_{\inf}-y_{0})\cdot\sum_{k=1}^{K}w_{k}*f_{k}(t,\Psi_{k})\\ \text{and \quad}p(\lambda;\theta)=\mathcal{E}=\{\sigma,y_{\inf}\}\cup\{w_{k}\}_{k=1}^{K}\cup\bigcup\limits_{k=1}^{K}\Psi_{k}\end{split}$ (6)
@@
$f_{k}(t,\Psi_{k})=f_{k}^{\prime}(x_{t},\Psi_{k})\text{\quad with \quad}x_{t}=\frac{1}{b_{max}}\cdot\begin{cases}t&\text{if }t\leq x_{sat,k}^{\lambda}\\ r_{sat,k}^{\lambda}(t-x_{sat,k})+x_{sat,k}&\text{if }t>x_{sat,k}^{\lambda}\end{cases}$ (7)
where $f_{k}$ has the following four parameters $\Psi_{k}$:
@@
Reference name Formula $f_{k}(x_{t},\Psi_{k})$ Prior $p(\Psi_{k})$
$pow\textsubscript{4}$ $((\epsilon_{1}^{\frac{-1}{\alpha_{1}}}-1)*\frac{x_{t}}{x_{sat,1}}+1)^{-\alpha_{1}}$ $\ln(\alpha_{1})\sim\mathcal{N}(1,1)$ $\log_{10}(x_{sat,k})\sim\mathcal{N}(0,1),\,\forall\,k$
$exp\textsubscript{4}$ $(\epsilon_{2})^{(\frac{x_{t}}{X_{sat,2}})^{\alpha_{2}}}$ $\ln(\alpha_{2})\sim\mathcal{N}(0,1)$ $\ln(\epsilon)\sim\mathcal{N}(-5,1),\,\forall\,k$
$ilog\textsubscript{4}$ $\frac{\ln(\alpha_{3})}{\ln((\alpha_{3}^{\frac{1}{\epsilon_{3}}}-\alpha_{3})\frac{x_{t}}{X_{sat,3}}+\alpha_{3})}$ $\ln(\alpha_{3}-1)\sim\mathcal{N}(-4,1)$ $y_{sat,k}=y_{\inf}-\epsilon\cdot(y_{\inf}-Y_{0}),\,\forall\,k$
$hill\textsubscript{4}$ $\frac{1}{(\frac{x_{t}}{X_{sat,4}})^{\alpha_{4}}(\frac{1}{\epsilon_{4}}-1)+1}$ $\ln(\alpha_{4})\sim\mathcal{N}(0.5,0.25)$ $1-r_{sat,k}\sim Exp(1),\,\forall\,k$
Table 2: The formulas for each of the four basis functions in our curve prior. Note that each of them are normalized to start at 1, converge to 0, and pass through the saturation point.
Finally, note that given these choices, we have $f_{comb}(t,\mathcal{E})\in[0,1]$ and we clip the Gaussian noise in $\hat{f}(t,\mathcal{E})$ in the same range. As a consequence, if performance does not naturally fall in this range, it must be normalized before passing it to FT-PFN. Examples of collections of curves generated using this prior can be found in Figure 5.
Figure 5: 21 i.i.d. samples of the FT-PFN prior, i.e., synthetically generated collections of learning curves for the same task using different hyperparameter configurations. In these examples, we consider 3 hyperparameters that are mapped onto the color of the curves, such that runs using similar hyperparameters, have similarly colored curves. We observe correlations, in varying degrees, between curves on the same task, especially with similar hyperparameter configurations.

### v2

@@ 4.1 Dynamic Surrogate Model (FT-PFN)
To sample a collection of curves on the same task from our prior data model, we (i) sample their hyperparameters from a unit hypercube; (ii) initialize $\operatorname*{\pi_{\text{config}}}$ as in (Müller et al., 2023); then (iii) apply Equation 2 to obtain the performance of each hyperparameter $\lambda$ at a given training time $t$. To reduce the entropy of this prior (and PFN training time), we assume the training time $t$ to be normalized in $[0,1]$. In practice, we also sample a maximum training time $\operatorname*{b_{\text{max}}}$ per task and define $t_{b}=\frac{b}{\operatorname*{b_{\text{max}}}}$ for $b\in[1,\operatorname*{b_{\text{max}}}]$. Further details about meta-training FT-PFN can be found in Appendix A, including the basis curves, their parameters, and illustrations of samples from our learning curve prior (Appendix A.1); a detailed description of the procedure we use for generating our meta-training data (Appendix A.2); and the architecture and hyperparameters used (Appendix A.3).
4.2 Dynamic Acquisition Policy (MFPI-random)
Following the Freeze-Thaw Bayesian Optimization framework (Section 3.2), we continue training the configuration that maximizes some acquisition function (AF), i.e., $\operatorname*{arg\!max}_{\lambda\in\Lambda}\texttt{AF}(\lambda)$. Conceptually, we would like to continue the training that is most likely to quickly produce a model performing substantially better than the best model obtained thus far. This notion can be formalized as an activation function,
$\texttt{MFPI}(\lambda;h,T)=\mathds{P}(\mathcal{M}{}(\lambda,~\min(b_{\lambda}+h,\operatorname*{b_{\text{max}}}))>T)$ (3)
which evaluates to the predicted likelihood that a candidate configuration $\lambda\in\Lambda$ after $h$ more steps of training obtains a model exceeding the $T$ performance threshold. Here, $\mathcal{M}{}$ is the trained surrogated model, and the extrapolation horizon $1\leq h\leq\operatorname*{b_{\text{max}}}$ is a parameter controlling the trade-off between immediate and long-term gains, i.e. what is quick enough, and the threshold $f_{\text{best}}\leq T<1$ a parameter controlling the trade-off between high and low risk/gain, i.e., what improvement is substantial. Note that for $h=\operatorname*{b_{\text{max}}}$ and $T=f_{\text{best}}$, we recover the Probability of Improvement (PI) acquisition function (Mockus et al., 1978). The values we choose for these hyper-hyperparameters will affect which configuration gets continued (see Figure 6, Appendix A.4). Their optimal settings depend on the desired freeze-thaw behavior and are not straightforward to determine. It might even be beneficial to adjust them dynamically during the run. Instead of using a fixed performance threshold $T$ or a fixed extrapolation horizon $h$ (Wistuba et al., 2022; Kadra et al., 2023), we explore a range of possible thresholds and horizons by randomizing these. Such random sampling procedure is undertaken every freeze-thaw BO iteration and is akin to an AF selection from a portfolio of different MFPIs. We posit that this hedging with a portfolio of AFs (portfolio of Equation 3 with different h,Th,~T) in each iteration benefits freeze-thaw setups where queries are more granular than standard BO. The result is a simple, parameter-free AF with a balanced exploration-exploitation trade-off,
@@ 4.2 Dynamic Acquisition Policy (MFPI-random)
$\begin{split}\texttt{MFPI-random}(\lambda)=\texttt{MFPI}(\lambda;h^{\text{rand}},~T^{\text{rand}})\\ \text{with }h^{\text{rand}}\sim\mathcal{U}(1,\operatorname*{b_{\text{max}}})\text{ and }\\ T^{\text{rand}}=f_{\text{best}}+\tau^{\text{rand}}\cdot(1-f_{\text{best}})\\ \text{ with }log_{10}(\tau^{\text{rand}})\sim\mathcal{U}(-4,-1)\end{split}$ (4)
Further details, as well as pseudo code (Algorithm 2) can be found in Appendix A.4.
@@ 5 Experiments
Our main baselines are both other recent freeze-thaw approaches: DyHPO (Wistuba et al., 2022) and DPL (Kadra et al., 2023). We reimplement the above two baselines in order to allow ablation of the online learning surrogates with different acquisition functions. Refer to Appendix C for more details.
@@ 5.1 Cost and Quality of Predictions
In this section, we compare the predictive capabilities of FT-PFN to that of existing surrogate models, including the deep Gaussian process of DyHPO (Wistuba et al., 2022) and the deep ensemble of power laws model of DPL (Kadra et al., 2023). We also consider a variant of FT-PFN trained on the same prior, but not taking the hyperparameters as input (referred to as “no HPs”). This variant bases its predictions solely on a set of partially observed learning curves.
@@
Figure 3: Comparison of our method against state-of-the-art baselines on all 3 benchmarks. First row shows normalized regret aggregated across multiple tasks in each benchmark (See Appendix B for benchmark details, and the results per task can be found in Appendix F). Second row shows the average ranks of each method.
@@ 5.2 Assessment of HPO performance
In this section, we present an extensive empirical comparison, showing the merits of our method on HPO tasks across a variety of tabular benchmarks (see, Appendix B) in the low budget regime. Additionally to DPL and DyHPO, we include Hyperband, ASHA, Gaussian process-based FT-BO (using DyHPO’s one-step EI acquisition function) and uniform random search, as baselines (see, Appendix C). Each algorithm is allocated a total budget of 1000 steps for every task, which corresponds to 20 full trainings on LCBench and Taskset  with each task being repeated 10 times with different seeds, each time starting from a different random configuration (see Algorithm 1). For PD1  tasks have varying maximum steps allowed ($\operatorname*{b_{\text{max}}}$), but for consistency and fair aggregation across benchmarks, we applied the same HPO budget of 1000 here too. We report two complementary metrics: The normalized regret, capturing performance differences, and the average rank of each method, capturing the relative order. Formally, the normalized regret corresponds to a $[0,1]$ normalization of the observed error w.r.t to the best (lowest) and worst (highest) errors recorded by all algorithms on the task.
@@
Figure 3 presents the comparative results per benchmark family. The results validate the superiority of freeze-thaw approaches (ifBO, DyHPO, and DPL) compared to standard approaches (Hyperband, ASHA, and random search) for low-budget settings. Most notably, these results establish the promise of ifBO, which either outperforms (on LCBench and Taskset) or competes closely (on PD1) with DPL and DyHPO. ifBO is also consistently the best on average rank across all benchmarks (see Appendix D.2, Figure 8). Appendix F (Figures 12-17) offers a closer look at the raw error metrics for each algorithm per task. These detailed results collectively confirm the robustness of ifBO for HPO tasks, showing its ability to compete if not outperform the most competitive baselines.
@@ 5.3 Ablation on Acquisition
In this section, we evaluate how ifBO performs in combination with other acquisition functions and aim to assess to what extent our novel acquisition function described in Section 4.2 contributes to its HPO success. To this end, we compare against ifBO variants combining FT-PFN with the EI-based acquisitions used in prior-art (Wistuba et al., 2022; Kadra et al., 2023), i.e., EI (one step) predicting one step in the future (DyHPO), and EI (max) predicting at the highest budget $b_{\text{max}}$ (DPL). We also include their PI counterparts PI (one step) and PI (max), as well as variants of MFPI-random that only vary the prediction horizon, PI (random horizon) with $T=0$, or only vary the threshold, PI (max, random-T) with $h=b_{\text{max}}$. Apart from the methods compared, the experimental setup is identical to that in Section 5.2.
@@
We perform comparable ablation studies for DPL and DyHPO, as detailed in Appendix D.3, to demonstrate the benefits of randomizing the horizon and the threshold.
@@ 6 Conclusion
Despite our promising results, we admit that attaining the sample efficiency required to scale up to modern deep learning (e.g., LLM pretraining) remains a challenging endeavor. Future work should attempt to extend our approach to take advantage of additional sources of prior information, e.g., to do in-context meta-learning, leveraging learning curves on related tasks (Ruhkopf et al., 2022); to incorporate user priors (Müller et al., 2023; Mallik et al., 2023a); and additional information about the training process (e.g., gradient statistics). An alternative to scaling up is scaling in parallel. Here, we expect our in-context learning approach to reduce the overhead further, as the online learning/refitting stage occurs on the critical path, while in-context learning during prediction can easily be parallelized. Finally, the current FT-PFN model has some limitations. First, it requires the performance metric and each hyperparameter value to be normalized in $[0,1]$; and supports up to 10 hyperparameters. While we believe that this is reasonable, future work building systems should push these limits, train larger models, on more data, and explore ways to scale to larger context sizes. In summary, there is a lot left unexplored, and we hope that the relative simplicity, efficiency, and public availability of our method, lowers the threshold for future research in this direction.
@@ Acknowledgements
We thank Johannes Hog for his feedback on an earlier version of the paper. Frank Hutter is a Hector Endowed Fellow at the ELLIS Institute Tübingen. All authors acknowledge funding by the state of Baden-Württemberg through bwHPC, the German Research Foundation (DFG) through grant numbers INST 39/963-1 FUGG and 417962828, and the European Union (via ERC Consolidator Grant Deep Learning 2.0, grant no. 101045765), TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ A.1 The Learning Curve Prior
$\operatorname*{\pi_{\text{curve}}}(\lambda,t)\sim\mathcal{N}\left(f_{\text{comb}}(t;\mathcal{E}),\sigma^{2}\right)$ $\text{with}\quad f_{\text{comb}}(t;\mathcal{E})=y_{0}+(y_{\infty}-y_{0})\cdot\sum_{k=1}^{K}w_{k}f_{k}(t;\Psi_{k})$
$\text{and}\quad(\sigma^{2},\underbrace{\left(y_{\infty},w_{1},\ldots,w_{K},\Psi_{1},\ldots,\Psi_{K}\right)}_{\mathcal{E}})\sim\operatorname*{\pi_{\text{config}}}(\lambda;\theta)$
Note that $\sigma^{2}$ and $\mathcal{E}$ are all outputs of the same neural network $\operatorname*{\pi_{\text{config}}}$. Due to the symmetry of this network, when marginalizing over $\lambda$ and $\theta$, all these parameters would have the same distribution. This is undesirable. To impose parameter-specific marginal distributions, we (i) estimate the empirical CDF of marginal output distribution; (ii) apply it to each output to obtain a new output with $\mathcal{U}(0,1)$ marginal distribution; (iii) apply the icdf of the parameter-specific target marginal distribution. Specifically, let $u_{1},u_{2},u_{3}\sim\mathcal{U}(0,1)$ be three i.i.d. uniform random variables that are hyperparameter independent and like $\theta$ are sampled once per task, then the non-basis curve specific parameters of our curve model are (marginally) distributed as follows:
$y_{\infty}\sim\mathcal{U}(y_{0},y_{max})\quad\text{with}\quad y_{0}=\min(u_{1},u_{2})\quad\text{and}\quad y_{max}=\begin{cases}\max(u_{1},u_{2})&\text{if }u_{3}\leq 0.25\\ 1.0&\text{if }u_{3}>0.25\end{cases}$
$\log(\sigma)\sim\mathcal{N}(-5,1)\quad\quad\quad\quad w_{k}=\frac{W_{k}}{W}\quad\text{with}\quad W_{k}\sim\mathrm{Gamma(1,1)}\quad\text{and}\quad W=\sum_{k=1}^{K}W_{k}$
@@
$f_{k}(t;\Psi_{k})=f_{k}^{\prime}(x_{t};\Psi_{k})\text{\quad with \quad}x_{t}=\begin{cases}t&\text{if }t\leq x_{\text{sat},k}^{\lambda}\\ r_{\text{sat},k}^{\lambda}(t-x_{\text{sat},k})+x_{\text{sat},k}&\text{if }t>x_{\text{sat},k}^{\lambda}\end{cases}$
where each $f_{k}$ has the following four parameters $\Psi_{k}$:
@@
Finally, note that given these choices, we have $f_{\text{comb}}(t,\mathcal{E})\in[0,1]$ and we clip the Gaussian noise in $\operatorname*{\pi_{\text{curve}}}(\lambda,t)$ in the same range. As a consequence, if performance does not naturally fall in this range, it must be normalized before passing it to FT-PFN. Examples of collections of curves generated using this prior can be found in Figure 5.
Reference name Formula $f^{\prime}_{k}(x_{t};\Psi_{k})$ Prior $p(\Psi_{k})$
pow4 $1-((\epsilon_{1}^{\frac{-1}{\alpha_{1}}}-1)*\frac{x_{t}}{x_{\text{sat},1}}+1)^{-\alpha_{1}}$ $\ln(\alpha_{1})\sim\mathcal{N}(1,1)$ $\log_{10}(x_{\text{sat},k})\sim\mathcal{N}(0,1),\,\forall\,k$
exp4 $1-(\epsilon_{2})^{(\frac{x_{t}}{X_{\text{sat},2}})^{\alpha_{2}}}$ $\ln(\alpha_{2})\sim\mathcal{N}(0,1)$ $\log_{1}0(\epsilon)\sim\mathcal{U}(-3,0),\,\forall\,k$
ilog4 $1-\frac{\ln(\alpha_{3})}{\ln((\alpha_{3}^{\frac{1}{\epsilon_{3}}}-\alpha_{3})\frac{x_{t}}{X_{\text{sat},3}}+\alpha_{3})}$ $\ln(\alpha_{3}-1)\sim\mathcal{N}(-4,1)$ $y_{\text{sat},k}=y_{\infty}-\epsilon\cdot(y_{\infty}-y_{0}),\,\forall\,k$
hill4 $1-\frac{1}{(\frac{x_{t}}{X_{\text{sat},4}})^{\alpha_{4}}(\frac{1}{\epsilon_{4}}-1)+1}$ $\ln(\alpha_{4})\sim\mathcal{N}(0.5,0.25)$ $1-r_{\text{sat},k}\sim Exp(1),\,\forall\,k$
Table 2: The formulas for each of the four basis functions in our curve prior. Note that each of them are normalized to start at 0, converge to 1, and pass through the saturation point.
Figure 5: Twenty-one i.i.d. samples of the FT-PFN prior, i.e., synthetically generated collections of learning curves for the same task using different hyperparameter configurations. In these examples, we consider 3 hyperparameters that are mapped onto the color of the curves, such that runs using similar hyperparameters, have similarly colored curves. We observe correlations, in varying degrees, between curves on the same task, especially with similar hyperparameter configurations.